Prove that the equation A+X=O has a unique solution in the set of all mxn matrices?

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

We have A+X =0
and also A+Y=0
Subtracting gives X-Y = 0 and so X = Y making X the unique solutioon

Regards